# HiRoPE: Length Extrapolation for Code Models

### v1

@@ Abstract
HiRoPE: Length Extrapolation for Code Models
@@ 2.1 Rotary Position Embedding in Transformer
Here $\mathbf{q}$ and $\mathbf{k}$ are the query and key vector for a specific attention head. At each layer, RoPE is applied on both query and key embeddings for computing attention scores. We can observe that the calculated attention score is only dependent on relative position $m-n$ through trigonometric functions, which reflects the core of RoPE that uses the absolute position to represent the relative distance. Existing studies show that when dealing with long plain text, the RoPE will meet O.O.D issues where the value of $m-n$ during inference is unseen Liu et al. (2023b), leading to poor performances.
@@ 3.1 Hierarchical format
$\begin{split}&\mathrm{hierarchicalAttn}=\mathrm{Re}\langle\mathbf{f}(\mathbf{q},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}}),\mathbf{f}(\mathbf{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}n_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{2}})\rangle\\ &=\sum_{i=0}^{d_{s}/2-1}[(q_{2i}k_{2i}+q_{2i+1}k_{2i+1})\cos(({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}})\theta_{i})\\ &\quad+(q_{2i}k_{2i+1}-q_{2i+1}k_{2i})\sin(({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}})\theta_{i})]\\ &\quad+\sum_{j=d_{s}/2+1}^{d/2-1}[(q_{2j}k_{2j}+q_{2j+1}k_{2j+1})\cos(({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\theta_{j})\\ &\quad+(q_{2j}k_{2j+1}-q_{2j+1}k_{2j})\sin(({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\theta_{j})]\\ &=:a\textquoteright({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\end{split}$ (4)
The attention score $a\textquoteright(\dots)$ ’ ( … ) we ultimately obtained through the inner product is quite elegant. It includes the relative position distance of various hierarchical levels, and this form can be similarly extended to the representation of more hierarchical positional structures. It indicates that the RoPE has the potential to be transformed into a hierarchical form, and we can use the hierarchical position index in the source code for this new form of RoPE, as shown in the left part of Figure 2.
@@ 4 Experiment Setup
RQ2. How is the language modeling capability of HiRoPE on long natural language sequences? The natural language lacks the explicit hierarchical structure information found in code, so we have made some modifications: we set every 128 tokens as a segment, and encode it as higher-level position information. We use the evaluation dataset from ReRoPE-eval Su (2023) in Section 5.2.
@@ 4.1 Base LLMs
Considering training on long context sequences is resource-intensive and time-consuming, we focus on those popular length extrapolation methods without training, including NTK bloc97 (2023), ReRoPE Su (2023) and Self-Extend Jin et al. (2024). These methods have shown impressive performance on long context language modeling.
@@ 5.1 Long Code Language Modeling
Table 3: Language Modeling Ability on CodeParrot-valid dataset. "nan" indicates that the model performs significantly poor on the given setting.
@@
Experiments show that original LLMs perform badly on the long code language modeling task. Even when the length slightly exceeds the pre-training length, the ppl of all models exceed 45, demonstrating their essential lack of modeling and understanding capabilities for longer codes. When we apply length extrapolation methods, all methods can reduce the loss and ppl into an acceptable range for those long source codes, and our HiRoPE almost outperforms all other baselines. Specifically, for ultra-long code sequences (length over 8192), HiRoPE achieves the best results in all metrics and under all settings. This fully demonstrates the advantages of our HiRoPE in modeling long sequence codes. Our method also shows generalization abilities. The four models evaluated have differences in model parameters, pre-training data, and pre-training length, yet our method has shown very good results on all these models.
@@ 7 Related Work
Existing large language models are originally trained with fixed context sizes. When dealing with longer sequences (such as long code), the model’s performance may decrease quite drastically. Recent studies have explored ways to expand the context length. For example, Position Interpolation Chen et al. (2023b) linearly down-scales the input position indices to match the original context window size of LLMs with several training steps. Similar studies Chen et al. (2023a); Peng et al. (2023); Chen et al. (2023c); Guo et al. (2023) also require fine-tuning. However, these methods all need additional tuning in longer contexts or face a disastrous collapse after the extrapolation bound. There are also some approaches without training. Some work use window attention to clip the long sequences such as Xiao et al. (2023); Han et al. (2023); Ding et al. (2023)). However, these methods rely on local information and may not effectively expand the context window, struggling with long dependencies in code. Recently, some methods have explored modifying the relative distance to extend the extrapolation length bloc97 (2023); Su (2023); Jin et al. (2024) and focus on the natural language text. We pursue the research line of training-free methods and propose considering the structural information of the code when modeling the position. We expand the traditional RoPE method into a hierarchical format and prove the effectiveness of our HiRoPE through theoretical derivations and practical experiments.
@@ 8 Conclusion
Limitation
@@
Appendix A Details of Code Symbol Understanding task
We extract long code files from popular open-sourced code repositories, especially those newly updated code projects to avoid data leakage. We construct a static analysis tool to get the abstract syntax tree of each code file, and then get all defined function and class names in it as the ground-truth output symbols. Details of these code projects are shown in Table 6. We show the number of files we extracted from each project as well as the average length of these code files. We also show the average number of symbols and their location statistics in the table.
ddbourgin/numpy-ml 2 14055 15 927 10845
gradio-app/gradio 1 12938 11 272 12265
huggingface/accelerate 4 12814.5 11.5 379 12500
huggingface/diffusers 4 13324.75 16.2 276 8347
huggingface/optimum 1 13491 11 580 12868
huggingface/peft 1 15457 11 514 8525
huggingface/transformers 18 12919.4 12.7 239 14514
langchain-ai/langchain/ 2 14195 11.5 285 10790
numpy/numpy 6 11158.8 14 100 12026
tensorflow/tensorflow 17 13191.7 13.4 285 14920
Total 56 12976.89 13.17 100 14920
Appendix B Details of ShortLLM experiments
@@
B.1 Theoretical calculation for O.O.D issues

### v2

@@ Abstract
HiRoPE: Length Extrapolation for Code Models Using Hierarchical Position
@@ 2.1 Rotary Position Embedding in Transformer
Here $\mathrm{Re}\langle\dots\rangle$ denotes the real part function for a complex number. $\mathbf{q}$ and $\mathbf{k}$ are the query and key vector for a specific attention head. At each layer, RoPE is applied on both query and key embeddings for computing attention scores. We can observe that the calculated attention score is only dependent on relative position $m-n$ through trigonometric functions, which reflects the core of RoPE that uses the absolute position to represent the relative distance. Existing studies show that when dealing with long plain text, the RoPE will meet O.O.D issues where the value of $m-n$ during inference is unseen Liu et al. (2023b), leading to poor performances.
@@ 3.1 Hierarchical format
$\begin{split}&\mathrm{hierarchicalAttn}=\mathrm{Re}\langle\mathbf{f}(\mathbf{q},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}}),\mathbf{f}(\mathbf{k},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}n_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}n_{2}})\rangle\\ &=\sum_{j=0}^{d_{s}/2-1}[(q_{2j}k_{2j}+q_{2j+1}k_{2j+1})\cos(({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}})\theta_{j})\\ &\quad+(q_{2j}k_{2j+1}-q_{2j+1}k_{2j})\sin(({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}})\theta_{j})]\\ &\quad+\sum_{j=d_{s}/2+1}^{d/2-1}[(q_{2j}k_{2j}+q_{2j+1}k_{2j+1})\cos(({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\theta_{j})\\ &\quad+(q_{2j}k_{2j+1}-q_{2j+1}k_{2j})\sin(({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\theta_{j})]\\ &=:a\textquoteright({\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}m_{1}-n_{1}},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}m_{2}-n_{2}})\end{split}$ (4)
The attention score $a\textquoteright(\dots)$ ’ ( … ) we ultimately obtained through the inner product is quite elegant. It includes the relative position distance of various hierarchical levels, and this form can be similarly extended to the representation of more hierarchical positional structures, as shown in Equation 5. Specifically, when $h=1$, it remains the same as the original RoPE.
$\begin{split}&\mathrm{hierarchicalAttn_{h}}\\ &=\mathrm{Re}\langle\mathbf{f}(\mathbf{q},m_{1},\dots,m_{h}),\mathbf{f}(\mathbf{k},n_{1},\dots,n_{h}))\rangle\\ &=:a\textquoteright(m_{1}-m_{2},\dots,m_{h}-n_{h})\end{split}$ (5)
It indicates that the RoPE has the potential to be transformed into a hierarchical form, and we can use the hierarchical position index in the source code for this new form of RoPE, as shown in the left part of Figure 2.
@@ 4 Experiment Setup
RQ2. How is the language modeling capability of HiRoPE on long natural language sequences? The natural language lacks the explicit hierarchical structure information found in code, so we have made some modifications: we set every 128 tokens as a segment, and encode it as higher-level position information. We use the evaluation dataset from ReRoPE-eval  [^1] https://github.com/bojone/rerope/blob/main/samples_15k.jsonl Su (2023) in Section 5.2. It is a dataset curated from Common Crawl Crawl (2023), refined by length-based selection criteria.
@@ 4.1 Base LLMs
4.2 Baselines
Considering training on long context sequences is resource-intensive and time-consuming, we focus on those popular length extrapolation methods without training, including NTK bloc97 (2023), ReRoPE Su (2023) and Self-Extend Jin et al. (2024). ReRoPE sets a window when calculating the relative distance, and clips those distances outside the window. Self-Extend uses group attention with the floor operation to calculate relative distances. These methods have shown impressive performance on long context language modeling. We also make comparisons with the original RoPE method, which is denoted as “origin” in the first row of Table 4, 5 and 6.
@@ 4.3 Inference Settings
4.4 Details of Code Symbol Understanding task
In RQ3, we design a new task: Code Symbol Understanding. Given a long code context, the model is required to output all the function names and class names defined in it. We extract long code files from popular open-sourced code repositories, especially those newly updated code projects to avoid data leakage. We construct a static analysis tool to get the abstract syntax tree of each code file, and then get all defined function and class names in it as the ground-truth output symbols. Details of these code projects are shown in Table 3. We show the number of files we extracted from each project as well as the average length of these code files. We also show the average number of symbols and their location statistics in the table.
ddbourgin/numpy-ml 2 14055 15 927 10845
gradio-app/gradio 1 12938 11 272 12265
huggingface/accelerate 4 12814.5 11.5 379 12500
huggingface/diffusers 4 13324.75 16.2 276 8347
huggingface/optimum 1 13491 11 580 12868
huggingface/peft 1 15457 11 514 8525
huggingface/transformers 18 12919.4 12.7 239 14514
langchain-ai/langchain/ 2 14195 11.5 285 10790
numpy/numpy 6 11158.8 14 100 12026
tensorflow/tensorflow 17 13191.7 13.4 285 14920
Total 56 12976.89 13.17 100 14920
@@ 5.1 Long Code Language Modeling
Table 4: Language Modeling Ability on CodeParrot-valid dataset. “nan” indicates that the model performs significantly poor on the given setting.
@@
Experiments show that original LLMs perform badly on the long code language modeling task. Even when the length slightly exceeds the pre-training length, the ppl of all models exceed 45, demonstrating their essential lack of modeling and understanding capabilities for longer codes. When we apply length extrapolation methods, all methods can reduce the loss and ppl into an acceptable range for those long source codes. Specifically, for ultra-long code sequences (length over 8192), HiRoPE achieves the best results in all metrics and under all settings. This fully demonstrates the advantages of our HiRoPE in modeling long sequence codes. Our method also shows generalization abilities. The four models evaluated have differences in model parameters, pre-training data, and pre-training length, yet our method has shown very good results on all these models.
@@ 7 Related Work
Existing large language models are originally trained with fixed context sizes. When dealing with longer sequences, the model’s performance may decrease quite drastically. Recent studies have explored ways to expand the context length. Position Interpolation Chen et al. (2023b) linearly down-scales the input position indices to match the original context window size of LLMs with several training steps. Similar studies Chen et al. (2023a); Peng et al. (2023); Chen et al. (2023c); Guo et al. (2023) also require fine-tuning. However, these methods need additional tuning in longer contexts or face a disastrous collapse after the extrapolation bound. There are also some approaches without training. Some studies use window attention to clip the long sequences such as Xiao et al. (2023); Han et al. (2023); Ding et al. (2023)). However, these methods rely on local information and may not effectively expand the context window, struggling with long dependencies in code. Recently, some methods have explored modifying the relative distance to extend the extrapolation length bloc97 (2023); Su (2023); Jin et al. (2024) and focus on the natural language text. We pursue the research line of training-free methods and propose considering the structural information of the code when modeling the position. Our work is inspired by the hierarchical embedding in our former work Zhang et al. (2023). We notice hierarchical positions are used in some concurrent work He et al. (2024); Lu et al. (2024) in NLP and CV domains. We expand the traditional RoPE method into a hierarchical format and prove the effectiveness of our HiRoPE through theoretical derivations and practical experiments.
@@ 8 Conclusion
9 Acknowledgments
This research is supported by the National Key R&D Program under Grant No. 2023YFB4503801, the National Natural Science Foundation of China under Grant No. 62072007, 62192733, 61832009, 62192730, and the Major Program (JD) of Hubei Province (No.2023BAA024).
Limitations
@@ 9 Acknowledgments
Appendix A Details of ShortLLM experiments
@@
A.1 Theoretical calculation for O.O.D issues